# Quality Time: Carbon-Aware Quality Adaptation for Energy-Intensive Services

### v2

@@ Abstract.
The energy demand of modern cloud services, particularly those related to generative AI, is increasing at an unprecedented pace. While hyperscalers collectively fail to meet their self-imposed emission reduction targets, they face increasing pressure from environmental sustainability reporting across many jurisdictions. To date, carbon-aware computing strategies have primarily focused on batch process scheduling or geo-distributed load balancing. However, such approaches are not applicable to services that require constant availability at specific locations due to latency, privacy, data, or infrastructure constraints.
In this paper, we explore how the carbon footprint of energy-intensive services can be reduced by adjusting the fraction of requests served by different service quality tiers. We show that adapting this quality of responses with respect to grid carbon intensity can lead to additional carbon savings beyond resource and energy efficiency. Building on this, we introduce a forecast-based multi-horizon optimization that reaches close-to-optimal carbon savings and is able to automatically adapt service quality for best-effort users to stay within an annual carbon budget. Our approach can reduce the emissions of large-scale LLM services, which we estimate at multiple 10,000 tons of CO2 annually, by up to 10 %.
@@
submissionid: ¡PAPER ID¿ccs: Social and professional topics Sustainabilityccs: Networks Cloud computingPreprint v2. Under review.
@@ 1. Introduction
Operating modern cloud services can require substantial computing resources and energy, which directly contributes to the rapidly growing carbon footprint of computing systems (S&P Globl Market Intelligence, 2023). Alongside traditional resource-intensive services like real-time data analytics, video processing, recommendation systems, and streaming, the growing use and high energy demand of generative AI is currently driving the expansion of data centers at a high pace (Bashir et al., 2024; Wu et al., 2024). As a result, major cloud providers, who collectively pledged to reduce their carbon footprint to zero by 2030, have been increasing their emissions significantly in recent years (Google, 2024; Microsoft, 2024; Meta Platforms, 2023) and are expected to row back on sustainability commitments (Lawrence, 2024).
Figure 1. The carbon intensity of electricity can vary significantly over time. Exploiting these temporal variations can reduce the carbon footprint of energy-intensive services.
Besides their social responsibility to minimize environmental impact, large companies are increasingly under pressure to report their greenhouse gas (GHG) emissions as part of environmental sustainability reporting (Project, 2024; initiative, SBTi). While reporting on Scope 2 emissions (GHG emissions caused by electricity consumption) is becoming mandatory in many parts of the world (Deloitte, 2024; European Parliament and Council of the European Union, 2022; UK Department for Business, Energy & Industrial Strategy, 2019; Japan Ministry of Enconomy, Trade and Industry (2023), METI), upcoming policy changes will mandate that large companies also disclose Scope 3 emissions (European Parliament and Council of the European Union, 2022; Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023)—including emissions from the use of cloud services. This will affect not only service providers, but also users of energy-intensive services.
@@
The concept of carbon-aware computing has emerged as a response to such challenges: Carbon-aware strategies aim to reduce emissions beyond energy efficiency by additionally taking into account how carbon-intensive the consumption of electricity at a current time and location is. Especially in the context of scheduling flexible batch workloads, the potential of temporal and spatial load shifting is well understood (Wiesner et al., 2021; Radovanovic et al., 2022; Lechowicz et al., 2023; Zheng et al., 2020; Wiesner et al., 2024; Hanafy et al., 2024b; Sukprasert et al., 2024b; Hanafy et al., 2024a). However, for continuously running services, where individual requests have little to no delay-tolerance, existing approaches exclusively focus on geographical load balancing (Gsteiger et al., 2024; Murillo et al., 2024; Souza et al., 2023; Zhou et al., 2016, 2013). However, many services cannot simply be distributed across regions due to data locality requirements, regulatory constraints, privacy and security concerns, or infrastructure limitations. For instance, generative AI inference typically relies on high-performance GPUs or specialized accelerators, which are often only available at specific locations. Carbon-aware approaches for services that are constrained to a single region remain largely unexplored.
@@
Figure 2. We investigate how adjusting the QoR with respect to carbon intensity can reduce emissions beyond savings related to reduced energy demand.
@@
In this work, we focus on the temporal variability of carbon intensity (see Figure 1) and explain how adapting service quality with respect to this metric does not only impact energy demand, but can yield additional carbon savings. We define the quality of responses (QoR) as a type of quality of service (QoS) metric that describes the proportion of requests served by different service quality tiers. For instance, machine learning models can be deployed in smaller, quantized, or pruned versions that use fewer resources and consume less energy, but compromise on response quality (Liang et al., 2021). Our approach minimizes the operational carbon emissions of energy-intensive services by optimizing resource provisioning and load balancing across service quality tiers, while ensuring compliance with service level objectives (SLOs) for different user groups.
@@
Figure 2 shows the accumulated energy usage and carbon emissions for an example scenario with two quality tiers over one week. Serving all requests at the high quality tier (QoR = 1) uses roughly twice the energy compared to serving them at the low quality tier (QoR = 0). Balancing requests evenly between the two tiers results in intermediate energy usage. We explore how strategically timing quality reductions based on carbon intensity can further reduce emissions by performing a comprehensive analysis across different regions, request patterns, and QoR constraints. Based on this, we propose an approach that achieves close-to-optimal savings despite real-life challenges such as forecast inaccuracies and the high computational complexity of the underlying optimization.
@@
    We formalize the problem of minimizing carbon emissions under QoR constraints through optimized resource provisioning and load balancing between service quality tiers.
@@
    We model emissions from the perspective of data center operators and cloud tenants (Scope 2 and 3 reporting) and prove that both models lead to equivalent decisions.
@@
    We propose a forecast-based multi-horizon optimization for minimizing service emissions, which can automatically adjust the QoR for best-effort users to stay within a predefined annual carbon budget.
  4. (4)
@@
    We evaluate the upper-bound potential and practicability of carbon-aware QoR adaptation by simulating a large-scale LLM inference service on eight real and synthetic request traces in ten different regions.
All experiments in this paper are fully reproducible. All code and data will be made publicly available upon acceptance.
2. Background and Motivation
Despite previous commitments to achieve fully carbon-free or even “carbon-negative” operations by 2030, Microsoft recently announced a 29.1 % rise in emissions since 2020 (Microsoft, 2024), Google reported a 48 % increase in GHG emissions within five years (Google, 2024), while Meta increased their emissions by 46 % in 2022 alone (Meta Platforms, 2023)—predominantly because of the training and deployment of generative AI. The rapid expansion of data centers, combined with the slow adoption of clean energy and efficiency measures, suggests that reaching net-zero emissions in the near future is highly unlikely (Lawrence, 2024).
All big hyperscalers have announced or introduced carbon reporting tools for their customers (Google Cloud, 2023; Microsoft, 2023; Amazon Web Services, 2023; Dixit and Tse, 2024). However, they are expected to significantly underreport emissions, due to opaque and inconsistent reporting methodologies (O’Brien, 2024). We anticipate an increasing focus on regulatory frameworks that will require cloud providers to disclose their carbon emissions more comprehensively.
Energy-intensive cloud services.  The rapid adoption of large language models (LLMs) requires immense processing power, not only for training but also for inference at scale (Faiz et al., 2024; Luccioni et al., 2024). As generative AI evolves towards multi-modality, emerging formats like video are expected to further increase future electricity demand (Li et al., 2024a).
Beyond generative AI, a range of current and emerging services are contributing to increasing energy demands. Video streaming, predominantly in high-resolution formats, now accounts for 82 % of all internet traffic (Systems, 2024). Platforms such as cloud gaming, virtual reality, and the metaverse require real-time graphics rendering, which consumes significant power (Stoll et al., 2022; Mills et al., 2019; Kattakinda et al., 2024). Additionally, scientific simulations such as climate modeling (Acosta et al., 2024), drug discovery (Wynendaele et al., 2021), astrophysics (Taffoni et al., 2019), fluid dynamics (Yang et al., 2024), and bioinformatics (Grealey et al., 2022) are growing in complexity and increasingly integrate AI services.
Growing pressure from GHG reporting.  The reporting of Scope 2 emissions, i.e. indirect GHG emissions caused by consuming electricity, is becoming increasingly standardized and mandatory across major jurisdictions (Deloitte, 2024; European Parliament and Council of the European Union, 2022; UK Department for Business, Energy & Industrial Strategy, 2019; Japan Ministry of Enconomy, Trade and Industry (2023), METI; Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023), with significant implications for data center operators, cloud tenants, and—ultimately—users of energy-intensive services.
In the US, the Securities and Exchange Commission (SEC) is advancing regulations that require public companies to fully disclose their Scope 2 emissions (Deloitte, 2024). While, the disclosure of Scope 3 emissions—i.e. all indirect emissions that occur across the value chain—stays voluntarily, for now, California soon requires large companies to also report Scope 3 emissions (Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023). Meanwhile, the EU’s Corporate Sustainability Reporting Directive (CSRD) mandates Scope 2 and 3 disclosures for large companies (European Parliament and Council of the European Union, 2022) beginning in 2024, aligning with the EU’s broader climate goals under the European Green Deal (European Commission, 2019). From 2026 on, this will also include non-EU companies with significant operations in the EU. As the significance of Scope 3 emissions has been emphasized repeatedly (Hettler and Graf‐Vlachy, 2024; Hertwich and Wood, 2018; Radonjič and Tompa, 2018), also jurisdictions outside the EU and California are likely to require Scope 3 emissions reporting in the future.
Implications for energy-intensive services.  When reporting emissions from the perspective of cloud service providers, we must distinguish between two cases:
  1. (1)
    They operate data centers themselves, in which case the operational carbon footprint is mainly determined by electricity usage, i.e. Scope 2 emissions.
  2. (2)
    They rent infrastructure from cloud providers, in which case the data center’s emissions will be passed down the value chain als Scope 3 emissions, which is an active topic of research (Amaral et al., 2023; Choochotkaew et al., 2023; Hè et al., 2023; Schmidt et al., 2023; Westerhof et al., 2023; Han et al., 2024).
Today’s, landscape is mixed: While, for example, Google relies on its proprietary infrastructure to train and serve AI models like Gemini (Pichai and Hassabis, 2024), Microsoft Azure provides the computational resources for operating OpenAI’s ChatGPT (OpenAI, 2023). We will address the two cases in our problem formalization, however, both are affected by stricter GHG reporting. Following, the increasing emissions of computing and the growing pressure for comprehensive accounting will soon have a direct impact on service cost. Carbon pricing mechanisms covered 23 % of global emissions in 2023, where revenues reached a record \$104 billion, and there is a clear trend towards stricter rules and higher prices (World Bank, 2024).
Assumptions on future regulatory policies.  As the effectiveness of annual and globally valid energy certificates is highly questionable (Bjørn et al., 2022; Holzapfel et al., 2023; Bashir et al., 2023), it is likely that future regulatory policies will demand time-based and location-based reporting of operational emissions. Based on current initiatives from industry (Energy, 2023; Cramer and Goodman, 2023) and legislation (Becker, 2021), we assume that operational carbon emissions of electricity usage will be reported hourly—or even sub-hourly (Haugen et al., 2024; Bashir et al., 2023).
We also follow the predominant assumption in related works that future regulations will use average carbon intensity (ACI), reported in carbon-equivalent GHG per unit of energy (gCO2/kWh), as a fundamental metric to determine the operational carbon emissions at specific times and locations (Gsteiger et al., 2024; Wiesner et al., 2021; Hanafy et al., 2024b; Sukprasert et al., 2024b; Radovanovic et al., 2022). ACI is defined as a weighted average of all produced energy in a region, weighted by the carbon intensity of the energy source. Some organizations, like Microsoft (Buchanan et al., 2023) advocate the use of marginal carbon intensity (MCI), which refers to the additional carbon emissions generated from meeting incremental demand at a given time. Although the two metrics can result in different scheduling decisions (Sukprasert et al., 2024a), both can be used in our framework. In our analysis, however, we assume ACI, as MCI is very ambiguous to compute in practice (Wiesner et al., 2021) and therefore a lot less likely to make it into official policies.
@@ 3. Carbon-Aware QoR Adaptation
Recently, there have been proposals to design applications that can intentionally sacrifice QoS, not only to manage economical interests or load peaks, but as a strategy to improve sustainability (Vitali, 2022; Vitali et al., 2023). Especially in AI inference, we have numerous options for fine-grained quality tiering that have different impacts on service quality and energy efficiency. For instance, service quality tiers for LLMs can vary based on factors such as the choice of model (e.g., number of parameters), quantization (ranging from 4-bit integers to 32-bit floating-point), and the number of tokens generated (i.e. providing shorter replies (Li et al., 2024b)), and can have a significant impact on the carbon footprint of inference.
In this work, we examine the impact of dynamically adjusting service QoR over time to enhance carbon efficiency within a single region, without relying on geo-distributed load balancing.
3.1. Problem Setting
We define an optimization time window $\left[0,T\right]$. Given an interval $\Delta$, we divide the window into $T/\Delta$ time intervals
(1) $\mathcal{T}_{i}=[i\Delta,(i+1)\Delta],\quad i=0,1,\dots,I-1,$
where $I=T/\Delta$ is the total number of intervals, i.e. $i=0$ corresponds to the first interval and $i=I-1$ to the last. For the remainder of this paper, we consider $T=1~{}\text{year}$ and $\Delta=1$ hour.
We define the set of user groups as $\mathcal{U}$, machine types $\mathcal{M}$ and service quality tiers $\mathcal{Q}$. Machine types can represent physical hardware as well as virtual machines (VMs) like cloud instances.
For each time interval, there is an associated carbon intensity $C^{i}$ and a number of requests $R^{i}$ per user group. Each machine can handle a certain number of requests by serving a certain quality tier (denoted as $K$), which corresponds to its current power usage $P^{i}$. The deployment $D^{i}$ must be sufficient to meet the number of incoming requests that are allocated to different service quality tiers ($A^{i}$) [^1] Note, that deployments can be constrained, i.e., $d^{i}_{m,q}\in[0,\dotsc,\bar{d}]$, where $\bar{d}$ represents the maximum number of available machines. Furthermore, when elements in $D^{i}$ are sufficiently large (i.e. scenarios with many active machines), an optimal result can be approximated by relaxing $D^{i}$ to continuous variables. This turns the optimization problem introduced later into a linear program that can be solved in polynomial time.. The key challenge lies is determining the optimal values for both $D^{i}$ and $A^{i}$ for each interval. Table 1 shows an overview of all input and decision variables.
Input variables
$\mathcal{U}$ Set of user groups $u$
$\mathcal{M}$ Set of machine types $m$
$\mathcal{Q}$ Set of service quality tiers $q$
$\gamma$ Validity period length
$R^{i}\in\mathbb{R}^{\left|\mathcal{U}\right|}$ Number of requests $\left(r^{i}_{u}\right)$ for user group $u$ during  $\mathcal{T}_{i}$.
$C^{i}\in\mathbb{R}$ Carbon intensity during $\mathcal{T}_{i}$.
$P^{i}\in\mathbb{R}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Power usage $\bigl{(}p^{i}_{m,q}\bigl{)}$ of machine $m$ serving quality $q$ during $\mathcal{T}_{i}$.
$K\in\mathbb{R}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Requests per timestep $\bigl{(}k_{m,q}\bigl{)}$ that machine $m$ can serve at quality $q$.
$C_{m}^{\text{emb}}$ Attributed embodied emissions for running machine $m$ for time $\Delta$.
Decision variables
$D^{i}\in\mathbb{N}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Deployment $\bigl{(}d^{i}_{m,q}\bigl{)}$, i.e., the number of machines $m$ running quality $q$ during $\mathcal{T}_{i}$.
$A^{i}\in\mathbb{R}^{\left|\mathcal{U}\right|\times\left|\mathcal{Q}\right|}$ Number of requests $\bigl{(}a^{i}_{u,q}\bigl{)}$ served by quality $q$ for user group $u$ during $\mathcal{T}_{i}$.
3.2. Quality of Responses
We define $\text{QoR}\in[0,1]$ as a metric that describes the proportion of requests served by different service quality tiers. In our subsequent analysis, we concentrate on a single user group, such as best-effort users, whose requests can be served by two distinct service tiers that differ in both response quality and energy demand. In this scenario, $\text{QoR}=0$ indicates that all requests are served by the lower-quality Tier 1, while $\text{QoR}=1$ indicates that all requests are served by the higher-quality but more energy-intensive Tier 2. When $\text{QoR}=0.5$, half of the requests are served by Tier 1 and half by Tier 2.
In the following, we introduce several metrics to establish a more general definition of QoR, that holds for multiple user groups and service quality tiers.
Service level objectives.  Adopting common terminology for quantifying QoS metrics (Keller and Ludwig, 2003), users of this framework can define a lower-bound service level objective $\underline{\text{SLO}}\in[0,1]^{\left|U\right|\times\left|Q\right|}$, where $\underline{\text{\text{slo}}}_{u,q}$ indicates the minimal proportion of requests that should be served by quality tier $q$ for user group $u$. Similarly, users can define an upper-bound $\overline{\text{SLO}}\in[0,1]^{\left|U\right|\times\left|Q\right|}$ to service quality. For example, LLM service providers often allow non-paying users a limited number of requests to access their highest-quality models, but do not intend to serve all requests at this quality tier.
@@ 3.2. Quality of Responses
$\underline{\text{SLO}}=\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\quad\text{and}\quad\overline{\text{SLO}}=\begin{bmatrix}0&1\\ 0.6&0.4\end{bmatrix}$
describes that premium users’ requests must always be served by Tier 2, while 60–100 % of requests by best-effort users can be served by Tier 1 and 0–40 % by Tier 2.
All requests per user group must be served by a quality tier:
(2) $\sum_{q\in\mathcal{Q}}\underline{\text{slo}}_{u,q}=1\quad\text{and}\quad\sum_{q\in\mathcal{Q}}\overline{\text{slo}}_{u,q}=1\quad\forall u\in\mathcal{U}.$
Service level indicators.  The service level indicator $\text{SLI}(\alpha,\omega)$ describes the observed distribution of requests across different quality tiers within the interval $\bigcup_{i=\alpha,\dotsc,\omega}\mathcal{T}_{i}$:
(3) $\text{sli}_{u,q}(\alpha,\omega)=\frac{\sum_{i=\alpha}^{\omega}a^{i}_{u,q}}{\sum_{i=\alpha}^{\omega}r^{i}_{u}},$
As the service level is constrained by the SLOs, each
(4) $\text{sli}_{u,q}(\alpha,\omega)\in\left[\min(\underline{\text{slo}}_{u,q},\overline{\text{slo}}_{u,q}),\max(\underline{\text{slo}}_{u,q},\overline{\text{slo}}_{u,q})\right].$
General definition of QoR.  We define the service’s quality of responses $\text{QoR}(\alpha,\omega)$ within an interval $\bigcup_{i=\alpha,\dotsc,\omega}\mathcal{T}_{i}$ through the relation between $\text{SLI}(\alpha,\omega)$, $\underline{\text{SLO}}$ and $\overline{\text{SLO}}$. In particular, for the set of parameters that permit optimization, defined as $\mathcal{V}=\{(u,q)\in\mathcal{U}\times\mathcal{Q}:\overline{\text{slo}}_{u,q}\neq\underline{\text{slo}}_{u,q}\}$, we define QoR as:
(5) $\text{QoR}(\alpha,\omega)=1-\min_{(u,q)\in\mathcal{V}}\left(\frac{\big{|}\overline{\text{slo}}_{u,q}-\text{sli}_{u,q}(\alpha,\omega)\big{|}}{\big{|}\overline{\text{slo}}_{u,q}-\underline{\text{slo}}_{u,q}\big{|}}\right).$
Following the above example, an $\text{SLI}(\alpha,\omega)=\begin{bmatrix}0&1\\ 0.7&0.3\end{bmatrix}$ would result in a $\text{QoR}(\alpha,\omega)=0.75$.
Validity periods.  As is typical in the assessment of QoS metrics (Keller and Ludwig, 2003), QoR is defined over validity periods $[\alpha,\alpha+1,...,\omega]$. These periods can either be disjoint (e.g., quarters, months, weeks) or—more commonly—overlapping. For the remainder of this paper, we assume that QoR is assessed over rolling windows of length $\gamma$ (integer multiple of adjustment interval $\Delta$). Staying with the above example, when the validity period is set to $\gamma=24~{}\text{h}$ and we want to guarantee a specific $\text{QoR}_{\text{target}}=0.5$, we have to ensure that within each 24 hour window at least 50 % of all requests are served by Tier 2. Formally, $0.5\leq\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=\alpha}^{\omega}$.
Figure 3 illustrates how the validity period length influences the hourly QoR over time. Longer validity periods provide greater flexibility in balancing service quality in response to fluctuations in carbon intensity, but can lead to prolonged periods of high or low hourly QoR.
@@
3.3. Power Model
To determine the operational carbon emissions from powering servers, we define the power usage $P^{i}=\bigl{(}p^{i}_{m,q}\bigl{)}$ of a machine $m$ running service quality $q$ during interval $\mathcal{T}_{i}$. We describe two power models from the perspective of a data center operator and a cloud tenant that report their Scope 2 and 3 emissions, respectively.
[…]
(6) $p^{i}_{m,q}=\text{PUE}\cdot\left(p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot\left(\text{util}^{i}_{q}\right)^{n}\right),$
where
(7) $\text{util}^{i}_{q}=\frac{\sum_{u\in\mathcal{U}}a^{i}_{u,q}}{\sum_{m\in\mathcal{M}}\left(d^{i}_{m,q}\cdot k_{m,q}\right)}.$
Cloud tenant perspective.  Service providers that rent their infrastructure from cloud providers, indirectly report the cloud data center’s power usage as part of their Scope 3 emissions. Although cloud providers recently started reporting emissions caused by renting VMs, their methods are inconsistent, lack transparency, and are subject to change. Until providers (are required to) adopt a standardized, publicly available methodology, we are left with estimating this factor. We assume that cloud providers attribute a constant, load-independent factor $P_{m}^{\text{attr}}$ to each active machine $m$:
(8) $p^{i}_{m,q}=p_{m}^{\text{attr}}$
3.4. Carbon Emission Model
We define a service’s estimated carbon emissions $E^{i}$ during time period $\mathcal{T}_{i}$ with respect to the region’s current operational carbon intensity $C^{i}$ (in gCO2/kWh) and the machine’s embodied emissions $C_{m}^{\text{ emb}}$ (in gCO2):
(9) $E^{i}=\sum_{m\in\mathcal{M}}\sum_{q\in\mathcal{Q}}d^{i}_{m,q}\cdot\left(p^{i}_{m,q}\cdot C^{i}+C_{m}^{\text{emb}}\right)$
Depending on methodology, embodied emissions might not always be reported with respect to infrastructure usage, in which case $C_{m}^{\text{emb}}=0$.
3.5. Optimization Problem
We formalize the problem of optimizing service deployments $D^{i}$ that minimize carbon emissions, while providing an allocation $A^{i}$ of requests to different service quality tiers that satisfies $\text{QoR}_{\text{target}}$:
@@ 3.5. Optimization Problem
where
  • All requests served (11) ensures that all requests during each interval $\mathcal{T}_{i}$ are attributed to a quality tier.
  • Sufficient resources (12) ensures that the provided machines have sufficient capacity to serve all requests per quality tier.
  • Respect SLOs (13) ensures that all SLOs are respected.
  • Sufficient QoR (14) ensures that the QoR of every validity period stays above $\text{QoR}_{\text{target}}$.
Note, that this model assumes knowledge of future carbon intensity ($C^{i}$) and requests ($R^{i}$) and can therefore not be solved in practice. However, we use it for assessing upper-bound potential and as the foundation of the online approach proposed in Section 4.
3.6. Complexity
The proposed optimization problem is an NP-hard mixed-integer linear program (MILP), as we minimize a linear objective function subject to linear constraints, which include discrete decision variables in $D^{i}$.
Equivalence of power models.  From the perspective of data center operators, the load-dependant power model from (6) introduces a non-linear relationship by multiplying values within $D^{i}$ and $A^{i}$. This would turn the MILP into an mixed-integer quadratic program, which is considerably harder to solve. However, during optimization, we can relax (6) to the same form as (8), as long as $n\leq 1$, i.e. the power model is concave.
Theorem 3.1.
We assume $p_{m,q}^{\text{max}}>p_{m}^{\text{idle}}>0$ and $n\in[0,1]$. When minimizing (10) subject to (11) – (14) from the perspective of data center operators, the utilization-dependent power model,
$p^{i}_{m,q}=p_{m}^{\text{idle}}+\bigl{(}p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\bigl{)}\cdot\bigl{(}\text{util}^{i}_{q}\bigl{)}^{n},$
can be simplified to
$p^{i}_{m,q}=p_{m,q}^{\text{max}},$
as it always yields an equivalent optimal deployment $D^{i}$.
The proof is provided in Appendix A. Intuitively, only if the power-law relationship is strictly convex ($n>1$), it can be more efficient to deploy multiple machines at lower utilization rather than fully utilizing fewer machines. However, this case cannot be observed in cloud data centers, where a main objective for energy efficiency is VM consolidation, to minimize the number of active machines (Xu et al., 2017; Gill and Buyya, 2018; Barroso and Hölzle, 2007). Hence, for $n\leq 1$, we can use a load-independent power model during optimization, as it yields the same optimal $D^{i}$.

### v3

# Carbon-Aware Quality Adaptation for Energy-Intensive Services

@@ Abstract.
The energy demand of modern cloud services, particularly those related to generative AI, is increasing at an unprecedented pace. To date, carbon-aware computing strategies have primarily focused on batch process scheduling or geo-distributed load balancing. However, such approaches are not applicable to services that require constant availability at specific locations due to latency, privacy, data, or infrastructure constraints.
In this paper, we explore how the carbon footprint of energy-intensive services can be reduced by adjusting the fraction of requests served by different service quality tiers. We show that adapting this quality of responses with respect to grid carbon intensity can lead to additional carbon savings beyond resource and energy efficiency and introduce a forecast-based multi-horizon optimization that reaches close-to-optimal carbon savings.
@@
journalyear: 2025copyright: ccconference: The 16th ACM International Conference on Future and Sustainable Energy Systems; June 17–20, 2025; Rotterdam, Netherlandsbooktitle: The 16th ACM International Conference on Future and Sustainable Energy Systems (E-ENERGY ’25), June 17–20, 2025, Rotterdam, Netherlandsdoi: 10.1145/3679240.3734614isbn: 979-8-4007-1125-1/2025/06ccs: Social and professional topics Sustainabilityccs: Networks Cloud computing
@@ 1. Introduction
Operating modern cloud services can require substantial computing resources and energy, which directly contributes to the rapidly growing carbon footprint of computing systems (S&P Globl Market Intelligence, 2023). Alongside traditional resource-intensive services like real-time data analytics, video processing, recommendation systems, and streaming, the growing use and high energy demand of generative AI is currently driving the expansion of data centers at a high pace (Bashir et al., 2024a; Wu et al., 2024). As a result, major cloud providers, who collectively pledged to reduce their carbon footprint to zero by 2030, have been increasing their emissions significantly in recent years (Google, 2024; Microsoft, 2024; Meta Platforms, 2023) and are expected to row back on sustainability commitments (Lawrence, 2024).
@@
The concept of carbon-aware computing has emerged as a response to such challenges: Carbon-aware strategies aim to reduce emissions beyond energy efficiency by additionally taking into account how carbon-intensive the consumption of electricity at a current time and location is. Especially in the context of scheduling flexible batch workloads, the potential of temporal and spatial load shifting is well understood (Wiesner et al., 2021; Radovanovic et al., 2022; Zheng et al., 2020; Wiesner et al., 2024; Hanafy et al., 2024b; Sukprasert et al., 2024; Hanafy et al., 2024a). In contrast, existing carbon-aware solutions for interactive services, which handle individual requests with minimal delay-tolerance, predominantly focus on geographical load balancing (Gsteiger et al., 2024; Murillo et al., 2024; Souza et al., 2023; Zhou et al., 2016, 2013). However, many services cannot simply be distributed across regions due to data locality requirements, regulatory constraints, privacy and security concerns, or infrastructure limitations. Carbon-aware approaches for services that are constrained to a single region remain largely unexplored.
@@
Figure 1. We investigate how adjusting the quality of responses (QoR) with respect to carbon intensity can reduce emissions beyond savings related to reduced energy demand.
@@
In this work, we focus on the temporal variability of carbon intensity and explain how adapting service quality with respect to this metric does not only impact energy demand, but can yield additional carbon savings. We define the quality of responses (QoR) as a type of quality of service (QoS) metric that describes the proportion of requests served by two different service quality tiers. For instance, the service quality and carbon footprint of LLM inference is determined by many factors such as the choice of model (Faiz et al., 2024), quantization and pruning (Liang et al., 2021), and the number of generated tokens (Li et al., 2024a).
@@
Figure 1 shows the accumulated energy usage and carbon emissions for an example scenario with two quality tiers over one week. Here, serving all requests at the high quality tier (QoR = 1) uses roughly twice the energy compared to serving them at the low quality tier (QoR = 0). Balancing requests evenly between the two tiers results in intermediate energy usage. We explore how strategically timing quality reductions based on carbon intensity can further reduce emissions by performing a comprehensive analysis across different regions, request patterns, and QoR constraints.
@@
    We formalize the problem of minimizing emissions under QoR constraints through optimized resource provisioning and load balancing between two service quality tiers.
@@
    We propose a forecast-based multi-horizon optimization for minimizing emissions that reaches close-to-optimal carbon savings.
@@
    We evaluate the potential and practicability of carbon-aware QoR adaptation by simulating a large-scale LLM service.
All code and data is publicly available [^1] https://github.com/dos-group/qualitytime.
@@ 2. Carbon-Aware QoR Adaptation
Recently, there have been proposals to design applications that can intentionally sacrifice QoS, not only to manage economical interests or load peaks, but as a strategy to improve sustainability (Vitali, 2022; Vitali et al., 2023). In this paper, we follow the predominant assumption that future regulations for Scope 2 emission reporting will be based on average carbon intensity (ACI), reported in carbon-equivalent emissions per unit of energy (gCO2/kWh) (Gsteiger et al., 2024; Wiesner et al., 2021; Hanafy et al., 2024b; Sukprasert et al., 2024; Radovanovic et al., 2022). We examine the impact of dynamically adjusting service QoR with respect to temporal changes in ACI to enhance carbon efficiency without relying on geo-distributed load balancing.
@@
Quality of Responses.  We define $\text{QoR}\in[0,1]$ as a metric that describes the proportion of requests served by two service quality tiers. That is, $\text{QoR}=0$ indicates that all requests are served by the lower-quality Tier 1, while $\text{QoR}=1$ indicates that all requests are served by the higher-quality but more energy-intensive Tier 2. When $\text{QoR}=0.5$, half of the requests are served by Tier 1 and half by Tier 2. Formally, for any interval $\bigcup_{i=\alpha,\dotsc,\omega}i$, QoR is defined as
(1) $\text{QoR}(\alpha,\omega)=\frac{\sum_{i=\alpha}^{\omega}a^{i}_{u,q}}{\sum_{i=\alpha}^{\omega}r^{i}_{u}},$
Validity periods.  As is typical in the assessment of QoS metrics (Keller and Ludwig, 2003), QoR is defined over validity periods of fixed length. These periods can either be disjoint (e.g., quarters, months, weeks) or, more commonly, overlapping. For the remainder of this paper, we assume that QoR is assessed over rolling windows of length $\gamma$ (integer multiple of adjustment interval $\Delta$). Staying with the above example, when the validity period is set to $\gamma=24~{}\text{h}$ and we want to guarantee a specific $\text{QoR}_{\text{target}}=0.5$, we have to ensure that within each 24 hour window at least 50 % of all requests are served by Tier 2. Formally, $0.5\leq\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=\alpha}^{\omega}$. Figure 2 illustrates how the validity period length influences the hourly QoR over time. Longer validity periods provide greater flexibility in balancing service quality in response to fluctuations in carbon intensity, but can lead to prolonged periods of high or low hourly QoR.
@@
Carbon emission model.  We define a service’s estimated carbon emissions $E^{i}$ during time interval $i$ with respect to the deployments current power usage, the region’s carbon intensity $C^{i}$ (in gCO2/kWh), and the attributed embodied emissions $C_{m}^{\text{ emb}}$ (in gCO2) for running machine $m$ for time $\Delta$:
(2) $E^{i}=\sum_{m\in\mathcal{M}}\sum_{q\in\mathcal{Q}}d^{i}_{m,q}\ \left(\Delta p^{i}_{m,q}\ C^{i}+C_{m}^{\text{emb}}\right)$
Depending on the methodology, embodied emissions may be excluded from scheduling decisions (Bashir et al., 2024b), in which case $C_{m}^{\text{emb}}=0$. We do not explicitly model emissions per user or request in this paper, but they can be derived using carbon attribution models (Han et al., 2024; Westerhof et al., 2023).
Optimization problem.  We formalize the problem of optimizing service deployments $D^{i}$ that minimize carbon emissions, while providing an allocation $A^{i}$ of requests to different service quality tiers that satisfies a $\text{QoR}_{\text{target}}\in[0,1]$:
@@
where (4) ensures that all requests during each interval $i$ are attributed to a quality tier, (5) ensures that the provided machines have sufficient capacity to serve all requests per quality tier, and (6) ensures that the QoR of every validity period stays above $\text{QoR}_{\text{target}}$.
Note, that this problem is not only NP-hard (see Appendix B) but also assumes knowledge of future carbon intensity ($C^{i}$) and requests ($R^{i}$). As such, it cannot be solved directly in practice.